A Bibliography of Publications on BMDP

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 January 2025 Version 1.11

Title word cross-reference

$10.50 [And79]. $16.20 [And79]. ^{(}R) [Ano84b]. P [And81, Mar78].

-Series [And81, Mar78].

0 [And79]. 0-520-03569-0 [And79].

1981 [Ano82]. 1985 [All86].

4V [DT79].

5V [KDV94a, HW91, KDV94b].

77 [And79, Hea78]. 77. [Mar78]. 79 [And81, Elt81, Hea82].

90 [Ano91, Eve91]. accompany [BMD89, BMD90, BMD92, Dix88, Dix90, Dix92]. ACO [SG80]. add [Ano90]. add-ons [Ano90]. additions [Dix83]. advanced [Chu83, Chu84, Chu85]. algorithm [WL93]. algorithms [Chu83, Chu84, Chu85]. alone [Ano89c]. analisis [SM82]. Analyse [Laf79, Laf81, Ber77]. analyses [SC87, Lac88, Zie88]. Analysis [DEL83, KDV94a, KDV94b, KW85, Pel95, Bro86, DT79, Mic70, Wit92, vdLVdL96]. Annotated [SG80]. Annotations [Bar83]. announces [Ano89b]. ANOVA [HP92]. applications [Chu83, Chu84, Chu85, Chu86, Chu87, Laf81]. Applied [SC87, Lac88, Zie88]. approaches [Fra78a, vdLVdL96]. Auswertungen [KW85]. available [Ano83b].

B [And79, And81, Elt81, Hea78, Hea82, Mar78].

BALANOVA [DEL83]. base [Chu86, Chu87]. Berkeley [And79, Uni83]. between [Fra79b]. Bibliography [Bee95]. biologiques [Laf79, Laf81]. Biomedical [And81, DB79, Elt81, Hea78, Hea82, Mar78, And79, Uni77, Dix64, Dix65, Dix67, Dix70b, Dix70a, Dix73, Dix75, DB77]. BMD [Uni77, Chu83, Chu84, Chu85, Chu86, Chu87, Dix64, Dix65, Dix67, Dix70b, Dix70a, Dix73, Fra76, Fra78a, Laf79, Laf81, Mic70, MBD84]. BMDP [KDV94a, Lac88, Zie88, And78, Ano83a, Ano84a, Ano84b, Ano89a, Ano89c, Ano89d, Ano89b, Ano90, Ano91, Uni83, Bar83, Bee95, BF78a, BF78b, Ber77, BD83, Bro86, DT79, DEL83, Dix75, DB77, Dix78, DB79, Dix81, Dix83, DB85, Dix88, Dix90, Dix92, DCT89a, DCT89b, EN88, EH92, Eve91, Fra76, Fra77, Fra78a, Fra78b, Fra79a, Fra79b, Fra81, FE82, HW91, Hil79, Hil82, Hil84, HP92, KDV94b, KW85, LG95, Laf79, Laf81, McC78, Mul78a, Mul78b, MBD84, Nel78, NFJN78, Pel95, PS02, Rya78, San78, SM82, Sch78, SG80, SC87, TF80, WL93, Wit92, And79, And81, Ano82, Cru84, Elt81, Hea78, Hea82, Mar78]. BMDP-5V [KDV94a, KDV94b]. BMDP-77 [DB77, Fra79b, And79, Hea78]. BMDP-77. [Mar78]. BMDP-79 [DB79, And81, Elt81, Hea82]. BMDP-V [SG80]. BMDP5V [vdLVdL96]. BMDP82 [Ano83b]. Book [And79, And81, Ano82, Cru84, Elt81, Hea78, Hea82, KDV94a, Lac88, Mar78, Zie88]. Brown [And79, Mar78, And81, Elt81, Hea78, Hea82].

C [And79]. California [And79]. Canonical [Mic70]. Categorical [Bro86]. CLISTs [TF80]. CMS [Uni83]. Comment [And78, Dix78, McC78, Nel78, NFJN78, Rya78, San78, Sch78]. Comparaison [Ber77]. Comparative [DEL83]. Comparing [vdLVdL96]. comparison [EN88, HP92, Pel95]. components [SG80].

Computer

[All86, And81, Bar83, Elt81, Fra76, Hea78, Hea82, Mar78, And79, Uni77, Dix64, Dix65, Dix67, Dix70b, Dix70a, Dix73, Dix75, DB77, DB79, Hil82, Hil84, SG80, All86]. computers [Ano83b]. condensed [Hil82, Hil84]. contingency [Mic70]. correlation [Mic70]. covariance [vdLVdL96]. Crunching [LG95].

data [Bro86, Chu86, Chu87, Fra78a, Fra78b, Pel95, vdLVdL96, KW85]. Database [Chu83, Chu84, Chu85]. Datenanalysessyteme [KW85]. datos [SM82]. del [SM82]. Design [Bar83]. designs [Wit92, vdLVdL96]. development [And79]. developments [Fra79b]. Difference [MBD84]. different [vdLVdL96]. digest [BMD89, BMD90, BMD92, Hil82, Hil84]. discriminant [WL93]. Distributed [Chu86, Chu87, HP92]. Dixon [And79, And81, Elt81, Hea78, Hea82, Mar78, Ano82, Cru84]. documentation [FE82]. d’ordinateur [Laf79].

Ed [Cru84]. edition [And79]. editor

[And79]. Einführung [KW85]. elementary [DCT89a, DCT89b]. Engelman [And79]. Engleman [Mar78]. estadisticos [SM82]. Examples [DEL83]. experiences [Wit92].

features [Fra79b]. February

[Chu83, Chu84]. file [Fra78b]. final

[Chu83, Chu84]. Five [KDV94a, KDV94b]. FORTRAN [FE82, Fra79a]. Frame [Mar78]. Frane [And79]. future [Bro86].

general [DT79]. GENMOD

[KDV94a, KDV94b]. GENSTAT [Ber77]. geograficos [SM82]. get [Uni83]. Getson [LG95]. growth [vdLVdL96]. guide [Fra79a, FE82, Hil82, Hil84].

Handbook [Lac88, Zie88, SC87]. hierarchical [vdLVdL96]. Highlights [Fra77]. HLM [KDV94a, KDV94b, vdLVdL96]. Hotel [All86].

IBM [Ano84a, TF80]. II [Fra79b]. incomplete [Wit92]. Input [Bar83]. interactively [TF80]. interest [Fra79b]. Interface [All86, Fra79b]. Introduccion [SM82]. introduces [Ano89c, Ano90, Ano91]. Introduction [KW85]. ISBN [And79].

J [And79, And81, Ano82, Cru84, Elt81, Hea78, Hea82, Lac88, Mar78, Zie88]. jack [WL93]. jack-knife [WL93]. January [Chu83, Chu84]. Jennrich [And79, Mar78].

Kentucky [All86]. knife [WL93].

L [And79, Mar78]. latest [Ano89a, Ano89b]. Lexington [All86]. linear [vdLVdL96]. logiciels [Ber77, Laf81]. Logistic [PS02]. longitudinal [vdLVdL96]. Ltd [EH92].

M [And79, And81, Elt81, Hea78, Hea82, Mar78]. major [Pel95]. management [Chu83, Chu84, Chu85, Chu86, Chu87]. manual [Dix88, Dix90, Dix92]. Manuals [And78, BF78a, BF78b, Dix78, McC78, Mul78a, Mul78b, Nel78, NFJN78, Rya78, San78, Sch78]. March [All86]. matrices [vdLVdL96]. matricielle [Laf79]. matrix [Fra78b]. measures [vdLVdL96]. microcomputer [Ano84b]. MINITAB [DCT89a, DCT89b, HP92, PS02]. mixed [vdLVdL96]. ML3 [KDV94a, KDV94b, vdLVdL96]. Modeling [PS02]. models [HP92, vdLVdL96]. Modern [DCT89a, DCT89b]. Multilevel [KDV94a, KDV94b]. Multiple [HP92]. multivariate [DT79]. multivariée [Laf79, Laf81]. MVS [TF80]. MVS/TSO [TF80].

non [Ano89c]. non-statisticians [Ano89c]. normally [HP92]. Novel [Mic70].

ons [Ano90]. Output [Bar83, SG80].

overview [Fra79a].

P

[Hea78, Hea82, DB77, DB79, And79, Elt81]. P-series [DB77, DB79, Elt81, And79]. package [Ano90]. Packages [Bar83, Pel95]. Pamela [LG95]. paquete [SM82]. Part [Fra79b]. PC [Ano84a, Ano89d, Ano91, EN88, Eve91]. PC-90 [Ano91, Eve91]. period [Chu83, Chu84]. Personal [LG95]. Plaza [All86]. pour [Ber77, Laf81]. pp [And79]. present [Bro86]. présentation [Laf79]. Press [And79]. prime [Ano83b]. printing [Dix83]. Pro [LG95]. probability [DCT89a, DCT89b]. procedures [HP92]. Proceedings [All86]. program [And79, Fra78b]. programas [SM82]. programmer [Fra79a, FE82]. programmes [Laf79]. programming [DCT89a, DCT89b]. Programs [And81, Elt81, Fra76, Hea78, Hea82, KDV94a, KDV94b, Mar78, And79, Ano89b, Uni77, BMD89, BMD90, BMD92, Dix64, Dix65, Dix67, Dix70b, Dix70a, Dix73, Dix75, DB77, DB79, Hil82, Hil84, Mic70]. Publications [Bee95].

quick [BMD89, BMD90, BMD92].

R [And79, Mar78]. Radisson [All86]. real [Chu86, Chu87]. real-time [Chu86, Chu87]. reference [BMD89, BMD90, BMD92]. Regression [PS02, vdLVdL96]. Rejoinder [BF78b, Mul78b]. release [Ano89a, Ano89d, Ano89b, BMD89, BMD90, BMD92, Dix88, Dix90, Dix92]. repeated [vdLVdL96]. report [Chu83, Chu84].

Research [Bar83]. Review [And79, And81, And78, Ano82, BF78a, BF78b, Cru84, Dix78, Elt81, Hea78, Hea82, KDV94a, KDV94b, Lac88, Mar78, McC78, Mul78a, Mul78b, Nel78, NFJN78, Rya78, San78, Sch78, Zie88, vdLVdL96, KDV94a]. Reviewed [LG95]. Reviews [EH92, HW91]. running [TF80].

Sample [Fra78b]. sampler [Hil79]. SAS [Bar83, DCT89a, DCT89b, EN88, Fra79b, HP92, KW85, Laf81, Pel95, PS02, Wit92]. science [All86]. Sciences [All86, Laf79, Laf81]. Series [Fra76, Fra77, And79, Dix70a, DB77, DB79, And81, Elt81, Hea78, Hea82, Mar78, And79]. Seventeenth [All86]. ships [Ano89d]. sigma [SG80]. Snell [Lac88, Zie88]. sociales [Laf79, Laf81]. Software [Ano82, EH92, HW91, Ano89a, Ano89c, Ano89d, Ano89b, Ano91, BD83, Dix81, Dix83, DB85, Dix88, Dix90, Dix92, HW91, Cru84]. SOLO [EH92, Ano89c, Ano90]. solution [Ano89c]. Some [Fra77, Fra79b]. source [FE82]. special [Fra79b]. specialisés [Ber77]. SPSS [And78, Bar83, BF78a, BF78b, Ber77, DEL83, Dix78, EN88, HP92, KW85, Laf79, Laf81, McC78, Mul78a, Mul78b, Nel78, NFJN78, Pel95, PS02, Rya78, San78, Sch78]. SPSSX [Bar83]. stand [Ano89c]. stand-alone [Ano89c]. started [Uni83]. STATA [PS02]. StatCat [Ano84b]. Statistical [Ano82, Ano91, Bar83, Cru84, EH92, Fra76, Fra77, HW91, KW85, Wit92, And79, Ano89a, Ano89c, Ano89d, Ano89b, BD83, Dix81, Dix83, DB85, Dix88, Dix90, Dix92, DCT89a, DCT89b, Pel95]. statisticians [Ano89c]. Statistics [Lac88, Zie88, All86, Ano90, DCT89a, DCT89b, SC87]. statistique [Ber77]. statistische [KW85]. stepwise [WL93]. Strategies [PS02]. structured [vdLVdL96]. subroutine [FE82]. supplement [Dix70a].

Survey [LG95]. survival [Pel95]. Symposium [All86]. Systat [PS02]. System [LG95]. Systems [KW85, And79].

tables [Mic70]. three [Pel95]. time

[Chu86, Chu87]. trois [Ber77]. TSO [TF80]. two [Ano89b, vdLVdL96].

unbalanced [Fra78a, Wit92]. univariate [DT79]. University [And79]. usage [Laf79]. user [BMD89, BMD90, BMD92, Hil82, Hil84]. users [Fra79b, WL93]. uses [Mic70]. using [vdLVdL96]. utilizacion [SM82].

V [SG80]. VARCL [KDV94b, KDV94a]. Variance [DEL83, DT79, SG80]. Version [EH92, Fra81, LG95].

W [And79, And81, Ano82, Cru84, Elt81, Hea78, Hea82, Mar78]. warning [WL93]. Windows [LG95]. write [Fra78b]. writeups [FE82].

X [Dix70a]. X-series [Dix70a]. xiii [And79].

References

Allen:1986:CSS

[All86] David M. Allen, editor. Computer science and statistics: Proceedings of the Seventeenth Symposium on the Interface of Computer Sciences and Statistics, Radisson Plaza Hotel, Lexington, Kentucky, March 17–19, 1985. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. URL http://

                                                   catalog.hathitrust.org/api/
                                                   volumes/oclc/13455449.html;

 http://www.dtic.mil/dtic/
tr/fulltext/u2/a166019.pdf;
 http://zbmath.org/?q=an:0647.
62008.

Andrews:1978:RMB

[And78] D. F. Andrews. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73 (361):85, March 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic).  URL http://www.
jstor.org/stable/2286521.

Anderson:1979:BRB

[And79] Gary D. Anderson. Book review:

BMDP-77, biomedical computer programs, P-series. by W. J. Dixon, series editor, and M. B. Brown editor 1977 edition. Systems, program and statistical development: L. Engelman, J. W. Frane and R. I. Jennrich. University of California Press, Berkeley, 1977. xiii + 880 pp. C $16.20, US $10.50. ISBN 0-520-03569-0. The Canadian Journal of Statistics = La revue canadienne de statistique, 7(1):111–113, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Anderson:1981:BRB

[And81] A. J. B. Anderson. Book review: BMDP-79: Biomedical Computer Programs P -Series, by W. J. Dixon, M. B. Brown. Biometrics, 37(1):204, March 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.
jstor.org/stable/2530547.

Anonymous:1982:BRBa

[Ano82]  Anonymous.  Book review:

BMDP Statistical Software 1981, by W. J. Dixon. Biometrics, 38(2):531, June 1982. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2530483.

Anonymous:1983:B

[Ano83a] Anonymous. BMDP. Computational Statistics & Data Analysis, 1(??):63–65, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0167947383900701.

Anonymous:1983:BAP

[Ano83b] Anonymous. BMDP82 available for prime computers. Computational Statistics & Data Analysis, 1(??):221, March 1983. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0167947383900932.

Anonymous:1984:BIP

[Ano84a] Anonymous. BMDP on the IBM PC. Computational Statistics & Data Analysis, 2(2): 180–182, August 1984. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0167947384900070.

Anonymous:1984:BSM

[Ano84b] Anonymous. The BMDP StatCat^{(}R) microcomputer. Computational Statistics & Data Analysis, 2(1):113–116, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0167947384900537.

Anonymous:1989:BSSa

[Ano89a] Anonymous. BMDP statistical software — latest release. Computational Statistics & Data Analysis, 7(3):323–325, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/016794738990039X.

Anonymous:1989:BSSd

[Ano89b] Anonymous. BMDP statistical software announces latest release with two new programs. Computational Statistics & Data Analysis, 8(2):218–219, July 1989. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0167947389900169.

Anonymous:1989:BSSb

[Ano89c] Anonymous. BMDP statistical software introduces Solo: the stand-alone solution for nonstatisticians. Computational Statistics & Data Analysis, 8 (2):214–216, July 1989. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0167947389900145.

Anonymous:1989:BSSc

[Ano89d] Anonymous. BMDP statistical software ships new PC release. Computational Statistics & Data Analysis, 8(2):216–217, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0167947389900157.

Anonymous:1990:BIN

[Ano90] Anonymous. BMDP introduces new add-ons for Solo statistics package. Computational Statistics & Data Analysis, 10 (1):95–98, August 1990. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/016794739090107S.

Anonymous:1991:SSB

[Ano91] Anonymous. Statistical software: BMDP introduces PC-90. Computational Statistics & Data Analysis, 11(2):227–230, March 1991. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/016794739190075D.

Barcikowski:1983:CPR

[Bar83] Robert S. Barcikowski. Computer Packages and Research Design: With Annotations of Input and Output from the BMDP, SAS, SPSS, and SPSSX Statistical Packages. University Press of America, Lanham, MD, USA, 1983. ISBN 0-8191-3494-5 (v. 1: paperback), 0-8191-3495-3 (v. 2: paperback), 0-8191-3496-1 (v. 3: paperback). xiv + 781 pp. LCCN QA276.4.C576 1983.

Brown:1983:BSS

[BD83] M. B. Brown and Wilfrid Joseph Dixon. BMDP statistical software. Berkeley, CA, USA, 1983 print. with additions edition, 1983. x + 733 pp. Includes index.

Beebe:1995:BPBa

[Bee95] Nelson H. F. Beebe. A bibliography of publications on BMDP. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1995. 9

pp. URL https://www.math.
utah.edu/pub/tex/bib/index-
table-b.html#bmdp. This report

is updated frequently.

Bernard:1977:CTL

[Ber77] G. Bernard. Comparaison de trois logiciels specialisés pour l’analyse statistique: GENSTAT, BMDP, SPSS. Technical report, Departement de mathématique, Université d’Orsay, Orsay, France, 1977. 15 pp.

Berk:1978:RMBa

[BF78a] Kenneth N. Berk and Ivor S.

Francis. A review of the manuals for BMDP and SPSS. Journal of the American Statistical Association, 73(361):65–71, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic).  URL http://www.
jstor.org/stable/2286517.

Berk:1978:RMBb

[BF78b] Kenneth N. Berk and Ivor S. Francis. A review of the manuals for BMDP and SPSS: Rejoinder. Journal of the American Statistical Association, 73 (361):94–95, March 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic).  URL http://www.
jstor.org/stable/2286527.

BMDP:1989:BUD

[BMD89] BMDP Statistical Software (Firm). BMDP user’s digest: quick reference for the BMDP programs: to accompany the 1988 software release. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1989. ISBN 0-935386-14-9. 218 pp. LCCN W 26.5 B661 1989.

BMDP:1990:BUD

[BMD90] BMDP Statistical Software (Firm). BMDP user’s digest: quick reference for the BMDP programs: to accompany the 1990 software release. Berkeley, CA, USA, 1990. 233 pp.

BMDP:1992:BUD

[BMD92] BMDP Statistical Software (Firm). BMDP user’s digest: quick reference for the BMDP programs: to accompany BMDP release 7. Berkeley, CA, USA, 1992. 251 pp. Developed by BMDP Statistical Software, Inc. Includes index.

Brown:1986:CDA

[Bro86] Morton B. Brown. Categorical data analysis in BMDP: present and future. In Allen [All86], pages 257–262. ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. US$65. URL

http://catalog.hathitrust.
org/api/volumes/oclc/13455449.
html;   http://www.dtic.mil/
dtic/tr/fulltext/u2/a166019.
pdf;   http://zbmath.org/?
q=an:0647.62008.

Chu:1983:DMA

[Chu83] Wesley W. Chu. Database management algorithms for advanced BMD applications: final report for the period from February 1, 1982 to January 31, 1983. Los Angeles, CA, USA, 1983. (various) pp. “Prepared for Ballistic Missile Defense Advanced Technology Center.” Includes bibliographies.

Chu:1984:DMA

[Chu84] Wesley W. Chu. Database management algorithms for advanced BMD applications: final report for the period from February 1, 1983 to January 31, 1984. Los Angeles, CA, USA, 1984. (various) pp. “Prepared for Ballistic Missile Defense Advanced Technology Center.” Includes bibliographies.

Chu:1985:DMA

[Chu85] Wesley W. Chu. Database management algorithms for advanced BMD applications. Los Angeles, CA, USA, 1985. 103 pp.

Chu:1986:DDB

[Chu86] Wesley W. Chu. Distributed data base management for real-time BMD applications. Los Angeles, CA, USA, 1986. (various) pp.

Chu:1987:DDB

[Chu87] Wesley W. Chu. Distributed data base management for real-time BMD applications. Los Angeles, CA, USA, 1987. (various) pp. Includes bibliographical references.

Crump:1984:BRB

[Cru84] Phelps P. Crump. Book review: BMDP Statistical Software (1981 Ed.) by W. J. Dixon. Technometrics, 26(4):416–417, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

                                        jstor.org/stable/1269510.

Dixon:1977:BBC

[DB77] Wilfrid Joseph Dixon and M. B. Brown. BMDP-77, biomedical computer programs, P-series. Berkeley, CA, USA, 1977. xiii + 880 pp. First published in 1975 under title: BMDP: biomedical computer programs.

Dixon:1979:BBC

[DB79] Wilfrid Joseph Dixon and M. B. Brown. BMDP-79: Biomedical computer programs P-series.

University of California Press, Berkeley, CA, USA, 2d print., rev. and updated edition, 1979. ISBN 0-520-03569-0. xiii + 880 pp. LCCN RA409.5 .C33 1979, Z699.5 M39 U6 1979. Includes index.

Dixon:1985:BSS

[DB85] Wilfrid Joseph Dixon and M. B.

Brown. BMDP statistical software. Berkeley, CA, USA, 1985 print edition, 1985. x + 733 pp. Includes index. Cover title: BMDP statistical software manual.

Dudewics:1989:MEP

[DCT89a] Edward J. Dudewics, Pinyuen Chen, and Baldeo K. Taneja. Modern elementary probability and statistics, with statistical programming in SAS, MINITAB, BMDP. American Sciences Press, Syracuse, NY, USA, 1989. ISBN 0-935950-19-2. ix + 375 pp. LCCN QA273 .D793 1989. US$34.50. Includes index. Includes bibliographical references.

Dudewicz:1989:MEP

[DCT89b] Edward J. Dudewicz, Pinyuen Chen, and Baldero Kumar Taneja. Modern elementary probability and statistics: with statistical programming in SAS, MINITAB, and BMDP. American Sciences Press, Syracuse, NY, USA, 1989. ISBN 0-935950-19-2. ix + 375 pp. LCCN QA273 .D793 1989. US$34.50. Includes index. Includes bibliographical references.

DeBoer:1983:AVC

[DEL83] Donald P. DeBoer, Frederick T. Evers, and Douglas Garfield Link. Analysis of Variance: Comparative Examples, BAL-ANOVA, BMDP and SPSS. Social Science Computing Laboratory, University of Western Ontario, London, Ontario, Canada, 1983. ISBN 0-7714-0427-1. (various) pp.

Dixon:1964:BBC

[Dix64] Wilfred Joseph Dixon. BMD: biomedical computer programs. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, 1964. vi + 585 pp. LCCN RA409.5 .C33 1964. Includes bibliographies.

Dixon:1965:BBC

[Dix65] Wilfrid Joseph Dixon. BMD: biomedical computer programs. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, rev edition, 1965. viii + 620 pp. LCCN RA409.5 .C33 1965. Work under grant FR-3 of the Division of Research Resource Facilities of the National Institutes of Health. Includes bibliographies.

Dixon:1967:BBC

[Dix67] Wilfred Joseph Dixon. BMD; biomedical computer programs. Berkeley, CA, USA, [2d ed.] edition, 1967. x + 600 pp. Includes bibliographies.

Dixon:1970:BBCb

[Dix70a] Wilfred Joseph Dixon. BMD; biomedical computer programs. X-series supplement. Berkeley, CA, USA, 1970. 260 pp.

Dixon:1970:BBCa

[Dix70b] Wilfrid Joseph Dixon. BMD; biomedical computer programs. University of California publications in automatic computation no. 2. University of California Press, Berkeley, CA, USA, [2d ed., rev.] edition, 1970. x + 600 pp. LCCN Z699.5 M39 U58 1970. Includes bibliographies.

Dixon:1973:BBC

[Dix73] Wilfrid Joseph Dixon. BMD; biomedical computer programs. University of California Press, Berkeley, CA, USA, third edition, 1973. ISBN 0-520-02426-5. vii + 773 pp. LCCN RA409.5 .C33 1973, Z699.5 M39 B13 1973. “This publication replaces numbers 2 and 3 of the series, Publications in automatic computation.” Includes bibliographical references.

Dixon:1975:BBC

[Dix75] Wilfrid Joseph Dixon. BMDP:

biomedical computer programs. University of California Press, Berkeley, CA, USA, 1975. xi + 792 pp. LCCN RA409.5.C33 1975. US$10.00. Edition of 1977 published under title: BMDP-77, biomedical computer programs, P-series.

Dixon:1978:RMB

[Dix78] W. J. Dixon. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73 (361):80–82, March 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/2286519.

Dixon:1981:BSS

[Dix81] Wilfrid Joseph Dixon. BMDP statistical software: 1981. Berkeley, CA, USA, 1981 ed edition, 1981. x + 725 pp. Includes index. On cover: Dept. of Biomathematics, University of California, Los Angeles.

Dixon:1983:BSS

[Dix83] Wilfrid Joseph Dixon. BMDP statistical software: 1983 printing with additions. University of California Press, Berkeley, CA, USA, 1983. ISBN 0-520-04408-8. x + 733 pp. LCCN RA409.5 .B57 1983. Includes index.

Dixon:1988:BSS

[Dix88] Wilfrid Joseph Dixon. BMDP statistical software manual: to accompany the 1988 software release. Berkeley, CA, USA, 1988. Includes index.

Dixon:1990:BSS

[Dix90] Wilfrid Joseph Dixon. BMDP statistical software manual: to accompany the 1990 software release. Berkeley, CA, USA, 1990. xviii + 1385 pp. Two volumes.

Includes index. Includes bibliographical references.

Dixon:1992:BSS

[Dix92] Wilfrid Joseph Dixon. BMDP statistical software manual: to accompany BMDP release 7. California/Princeton Fulfillment Services [distributor], Princeton, NJ, USA, 1992. ISBN 0-520-08141-2 (set), 0-520-08138-2 (v. 1), 0-520-08139-0 (v. 2), 0-520-08140-4 (v. 3). (various) pp. LCCN QA276.4 .B58 1992. US$72.50 (set), US$33.00 (v. 1), US$33.00 (v. 2), US$22.00 (v. 3). Volume 3 has a distinctive title: The data manager. Includes bibliographical references and indexes.

Davidson:1979:BUM

[DT79] Michael L. Davidson and Jerome D.

Toporek. BMDP 4V–general univariate and multivariate analysis of variance. Technical report / BMDP Statistical Software; no. 67. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, rev edition, 1979. 117 pp. LCCN CM.2.2.05. Includes bibliography.

Evans:1992:SSR

[EH92] Stephen Evans and Martin Hurley. Statistical software reviews: SOLO: Version 3.0, by BMDP Statistical Software Ltd. Applied Statistics, 41(3):605–609, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Elton:1981:BRB

[Elt81] R. A. Elton. Book review: BMDP-79: Biomedical Computer Programs, P-Series, by W. J. Dixon; M. B. Brown. Journal of the Royal Statistical Society. Series A (General), 144 (2):263–264, ???? 1981. CO-DEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.
                                        org/stable/2981928.

Evans:1988:CSP

[EN88] Gerald Evans and William A. Newman. A comparison of SPSS PC, SAS PC, and BMDP PC. Collegiate Microcomputer, 6(2): 97–106, May 1988. CODEN CMICDL. ISSN 0731-4213.

Everitt:1991:BP

[Eve91] B. S. Everitt. BMDP PC-90. Statistics and Computing, 1(1):71–73, September 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF01890838.

Frane:1982:BPG

[FE82] James W. Frane and Laszlo Engelman. BMDP programmer’s guide and subroutine writeups: documentation of the FORTRAN source. Bmdp technical report; no. 55, BMDP Statistical Software, Dept. of Biomathematics, University of California, Los Angeles, CA, USA, 1982. 449 pp.

Frane:1976:BBS

[Fra76] J. W. Frane. The BMD and BMDP series of statistical computer programs. Communications of the ACM, 19:570– 576, October 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frane:1977:SSH

[Fra77] James W. Frane. Some statistical highlights of the BMDP series. The American Statistician, 31(3):131–133, August 1977. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL http://www.
jstor.org/stable/2682965.

Frane:1978:BBA

[Fra78a] James W. Frane. BMD and BMDP approaches to unbalanced data. Technical report / Health Sciences Computing Facility, University of California, Los Angeles; no. 41. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, 1978. 10 pp. LCCN CM.2.2.02. Includes bibliography.

Frane:1978:SPW

[Fra78b] James W. Frane. Sample program to write a data matrix on a BMDP file. Technical report / Health Sciences Computing Facility, University of California, Los Angeles; no. 42. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, 1978. 5 pp. LCCN CM.2.2.02. Includes bibliography.

Frane:1979:BFP

[Fra79a] James W. Frane. The BMDP Fortran programmer’s guide: an overview. Technical report / BMDP Statistical Software; no.

66. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1979. 6 pp. LCCN CM.2.2.05.

Frane:1979:PII

[Fra79b] James W. Frane. Part I–The 1978 interface between BMDP-77 and SAS and some BMDP features of interest to SAS users. Part II–Recent developments in BMDP of special interest to SAS users. Technical report / BMDP Statistical Software; no.

40. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1979. 4 + 2 + 8 pp. LCCN CM.2.2.02. Includes bibliography.

Frane:1981:VB

[Fra81] James W. Frane. The 1981 version of BMDP. The American Statistician, 35(4):261– 262, November 1981. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/
stable/2683306.

Hearne:1978:BRB

[Hea78] F. Terry Hearne. Book review:

BMDP-77 Biomedical Computer Programs, P Series by W. J. Dixon; M. B. Brown. Technometrics, 20(3):320, August 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic).  URL http://www.
jstor.org/stable/1268144.

Hearne:1982:BRB

[Hea82] F. Terry Hearne. Book review:

BMDP-79 Biomedical Computer Programs, P Series by W. J. Dixon; M. B. Brown. Technometrics, 24(1):82, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic).  URL http://www.
jstor.org/stable/1267584.

Hill:1979:BS

[Hil79] Mary Ann Hill. BMDP sampler.

Technical report / BMDP Statistical Software; no. 53. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1979. 93 pp. LCCN CM.2.2.03. Includes “Complete list of BMDP programs.”.

Hill:1982:BUD

[Hil82] Mary Ann Hill. BMDP user’s digest: a condensed guide to the BMDP computer programs. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 2nd ed edition, 1982. ISBN 0-935386-02-5. 157 pp. LCCN W 26.5 H647b 1982.

Hill:1984:BUD

[Hil84] Mary Ann Hill. BMDP user’s digest: a condensed guide to the BMDP computer programs. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 3rd ed edition, 1984. 177 pp. LCCN QA276.4 .H54 1984, W 26.5 H647b 1984.

Holzer:1992:MCP

[HP92] Christian Holzer and Manfred Precht. Multiple comparison procedures for normally distributed ANOVA models in SAS, SPSS, BMDP, and MINITAB. Computational Statistics & Data Analysis, 13(3):351–358, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0167947392901412.

Hand:1991:SSR

[HW91] David J. Hand and Maeve Wall. Statistical software reviews: BMDP 5V, by BMDP Statistical Software. Applied Statistics, 40(1):192–194, 1991. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kreft:1994:BRB

[KDV94a] Ita G. G. Kreft, Jan De Leeuw, and Rien Van Der Leeden. Book review: Review of Five Multilevel Analysis Programs: BMDP-5V, GENMOD, HLM, ML3, VARCL. The American Statistician, 48(4):

324–335, November 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL http://www.
jstor.org/stable/2684845.

Kreft:1994:RFM

[KDV94b] Ita G. G. Kreft, Jan De Leeuw, and Rien Van Der Leeden. Review of five multilevel analysis programs: BMDP-5V, GEN-MOD, HLM, ML3, and VARCL. The American Statistician, 48(4): 324–??, November 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kuffner:1985:DSA

[KW85] Helmuth Kuffner and Reinhard Wittenberg. Datenanalysessyteme fur statistische Auswertungen: e. Einführung in SPSS, BMDP und SAS [English: Data Analysis Systems for Statistical ??: An Introduction to SPSS, BMDP, and SAS]. G. Fischer, Stuttgart, Germany, 1985. 292 pp.

Lachenbruch:1988:BRA

[Lac88] P. A. Lachenbruch. Book review: Applied Statistics: a Handbook of BMDP Analyses, by E. J. Snell. Biometrics, 44(3): 918–919, September 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/
stable/2531614.

Laforge:1979:AMS

[Laf79] Hubert Laforge. Analyse multivariée: en sciences sociales et biologiques: présentation matricielle et usage des programmes d’ordinateur BMD, BMDP et SPSS. Presses de l’Université Laval, Québec, Québec, Canada, 1979. vi + 450 pp.

Laforge:1981:AMP

[Laf81] Hubert Laforge. Analyse multivariée: pour les sciences sociales et biologiques avec applications des logiciels BMD, BMDP, SPSS, SAS. Etudes vivantes, Saint-Lambert, Québec, 1981. ISBN 2-7607-0075-5. xiii + 317 pp.

Lachenbruch:1995:PCB

[LG95] Peter A. Lachenbruch and Pamela Getson. Personal crunching: BMDP new system for Windows, version 1.11, reviewed by Pamela Getson, and Survey Pro for Windows, version 1.8. Chance, 8(4):53–59, 1995. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Marriott:1978:BRBc

[Mar78] F. H. C. Marriott. Book review: BMDP-77. Biomedical Computer Programs, P -Series, by W. J. Dixon; M. B. Brown; L. Engleman; J. W. Frame; R. I. Jennrich. Journal of the Royal Statistical Society. Series A (General), 141(3):411, ???? 1978. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.
                                        org/stable/2344816.

Musser-Brauner:1984:BBW

[MBD84] Melonia L. Musser-Brauner and W. J. Dixon. BMD or BMDP? What’s the difference? The American Statistician, 38(4): 318–319, November 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/
stable/2683421.

McCarthy:1978:RMB

[McC78] John L. McCarthy. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361):86–89, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic).  URL http://www.
jstor.org/stable/2286522.

Mickey:1970:NUB

[Mic70]  M. R. Mickey.  Novel uses

of BMD programs. I. Canonical correlation analysis of contingency tables. Technical report / Health Sciences Computing Facility, University of California, Los Angeles; no. 2. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, 1970. 8 pp. LCCN CM.2.2.01. Includes bibliography.

Muller:1978:RMBa

[Mul78a] Mervin E. Muller. A review of the manuals for BMDP and SPSS. Journal of the American Statistical Association, 73 (361):71–80, March 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/2286518.

Muller:1978:RMBb

[Mul78b] Mervin E. Muller. A review of the manuals for BMDP and SPSS: Rejoinder. Journal of the American Statistical Association, 73(361):95–98, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/2286528.

Nelder:1978:RMB

[Nel78] J. A. Nelder. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73 (361):89–90, March 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/2286523.

Nie:1978:RMB

[NFJN78] Norman H. Nie, Jonathan B. Fry, Jean G. Jenkins, and Marija J. Norusis. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73 (361):82–84, March 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/2286520.

Pelz:1995:ASD

[Pel95] Corey J. Pelz. Analysis of survival data: a comparison of three major statistical packages (SAS, SPSS, BMDP). Thesis, (m.s.), Medical College of Wisconsin, ????, 1995. 128 pp.

Peng:2002:MSL

[PS02] Chao-Ying Joanne Peng and Tak-Shing Harry So. Modeling strategies in logistic regression with SAS, SPSS, Systat, BMDP, Minitab, and STATA. Journal of Modern Applied Statistical Methods, 1(1):??, ???? 2002. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/
article/view/22.

Ryan:1978:RMB

[Rya78] Thomas A. Ryan, Jr. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361):90–91, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic).  URL http://www.
jstor.org/stable/2286524.

Sande:1978:RMB

[San78] Gordon Sande. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73 (361):91–92, March 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic).  URL http://www.
jstor.org/stable/2286525.

Snell:1987:ASH

[SC87] E. J. Snell and D. R. Cox. Applied statistics: a handbook of BMDP analyses. Chapman and Hall, Ltd., London, UK, 1987. ix + 171 pp. LCCN QA276 .C66531 1981 Suppl. “A complement to Applied statistics: principles and examples [by] D. R. Cox and E. J. Snell.” Includes index. On t.p. the registered trademark symbol “TM” is superscript following “BMDP” in the title.

Schucany:1978:RMB

[Sch78] William R. Schucany. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73(361):92–93, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/2286526.

Searle:1980:ACO

[SG80] Shayle R. Searle and B. A. Grimes. Annotated computer output for variance components (ACO [sigma]2): BMDP-V. Paper … in the Biometrics Unit; no. BU-716-M. Biometrics Unit, Cornell University, 1980. xxiv + 45 pp.

Santos-Preciado:1982:IUP

[SM82] Jose Miguel Santos Preciado and Carmen Muguruza Canas. Introduccion, a la utilizacion del paquete de programas estadisticos BMDP en el analisis de datos geograficos. Departamento de Geografia Humana, Facultad de Geografia e Historia, Universidad Complutense, Madrid, Spain, 1982. ISBN 84-7491-051-X. 189 pp. LCCN G70.2 .S26 1982.

Toporek:1980:IMC

[TF80] Jerome D. Toporek and James W.

Frane. IBM MVS/TSO CLISTs for running BMDP interactively. Technical report / BMDP Statistical Software; no. 78. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1980. 11 pp.

BMDP:1977:BBC

[Uni77] University of California, Los Angeles.Health Sciences Computing Facility. BMD, biomedical computer programs. University of California Press, Berkeley, CA, USA, 1977. 67–764 pp. Includes bibliographies.

BMDP:1983:BUC

[Uni83] University of California, Berkeley.Computing Services. BMDP under CMS: how to get started with BMDP at Berkeley. Computing Services, University of California, 1983. 12 pp. LCCN CM.3.2.01. Includes bibliography.

vanderLeeden:1996:RTD

[vdLVdL96] Rien van der Leeden, Karen Vrijburg, and Jan de Leeuw. A review of two different approaches for the analysis of growth data using longitudinal mixed linear models: Comparing hierarchical linear regression (ML3, HLM) and repeated measures designs with structured covariance matrices (BMDP5V). Computational Statistics & Data Analysis, 21(5):583–605, May 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0167947396822961.

Wittkowski:1992:SAU

[Wit92] Knut M. Wittkowski. Statistical analysis of unbalanced and incomplete designs — experiences with BMDP and SAS. Computational Statistics & Data Analysis, 14(1):119–124, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/016794739290086U.

White:1993:JKS

[WL93] A. P. White and W. Z. Liu. The jack-knife with a stepwise discriminant algorithm — a warning to BMDP users. Journal of Applied Statistics, 20(1):187–190, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ziegel:1988:BRBaa

[Zie88] Eric Ziegel. Book review: Applied Statistics: A Handbook of BMDP Analyses by E. J. Snell. Technometrics, 30(4): 468, November 1988. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic).  URL http://www.
jstor.org/stable/1269830.